# Reconstruction of Quantum Fields

### v1

@@ Abstract
One of the traditional ways of introducing bosons and fermions is through creation and annihilation algebras. Historically, these have been associated with emission and absorption processes at the quantum level and are characteristic of the language of second quantization. In this work, we formulate the transition from first to second quantization by taking quotients of the state spaces of distinguishable particles, so that the resulting equivalence classes identify states that contain no information capable of distinguishing between particles, thereby generalising the usual symmetrisation procedure. Assuming that the resulting indistinguishable-particle space (i) admits an ordered basis compatible with how an observer may label the accessible modes, (ii) is invariant under unitary transformations of those modes, and (iii) supports particle counting as a mode-wise local operation, we derive a new class of creation–annihilation algebras. These algebras reproduce the partition functions of transtatistics–maximal generalisations of bosons and fermions consistent with these operational principles.
@@ I Introduction
which describes distinguishable particles whose single–particle spaces are isomorphic. Indistinguishability is then imposed by projecting onto the symmetric or antisymmetric subspaces, yielding the usual occupation–number description in which only the number of particles per accessible mode is relevant. In this traditional formalism the exclusivity of Bose and Fermi statistics is assumed rather than derived, and any alternative statistical behaviour is excluded a priori. Nonetheless, a variety of classical and recent works have exhibited consistent generalisations, including parastatistics Green53; Greenberg1964; HARTLE69; ryan1963; stolt1970; stoilova, quons quons; Greenberg1991, Gentile statistics Gentile40, and several topological or combinatorial extensions dewitt; myrheim; Wilczek82; Read_2003; fredenhagen89; Surya_2004; Balachandran_2001; Baez06; vanenk2019; ZhangHanssonKivelson1989; Witten1989; Haldane08; Laughlin1983. These are typically introduced by ad hoc modifications of commutation relations, rather than through a structural analysis of indistinguishability.
In this work we develop a systematic treatment of the passage from distinguishable to indistinguishable particles based on operational constraints. We take as starting point again the algebra $T(H)$ whose elements represent labelled multi–particle states, following Segal’s tensor-algebra formulation of second quantization Segal1956a; Segal1956b; Segal1958. Indistinguishability is implemented by an equivalence relation that removes precisely the information operationally inaccessible for distinguishing one constituent from another. The quotient
@@
To connect this with experimental structure, we assume a factorisation
@@ II.1 Operational indistinguishability and information loss
Consider a family of systems that, at the single–particle level, are described by the same Hilbert space $H$. An individual particle state is represented by a unit vector $\psi\in H$, and an collection of $n$ labelled particles is described by a state in $H^{\otimes n}$. The labels encode individuality: the factor $H$ in each tensor position corresponds to a specific system, and permutations of the factors implement relabellings.
@@
In the standard approach, indistinguishability is represented by the action of the symmetric group $S_{n}$ on $H^{\otimes n}$, and the subspace of physical states is obtained by projecting onto the totally symmetric or totally antisymmetric subspace. This can be rephrased as choosing $I$ to be the ideal generated by the antisymmetric or symmetric components in degree two, respectively. The equivalence relation introduced in that case is the one to corresponds to the identification of states that only differ on a permutation of the labels. Depending on how the group $S_{n}$ is acting on the states, we will get either the symmetric or antisymmetric states.
@@ II.2 Fock space as a quotient algebra:
carries a natural $\mathbb{N}_{0}$–grading by tensor degree, what determines the number of particles. We assume that the ideal $I$ is homogeneous,
@@
In this way the transition from distinguishable to indistinguishable particles does not change the number of particles, as we can expect. Additionally we see that the degree–zero component $F_{0}$ is one–dimensional, spanned by the vacuum $|0\rangle$, and the degree–one component $F_{1}$ is naturally isomorphic to the space of 1-particle $H$.
@@ II.3 Segal’s theorem
The construction we employ was used by Segal where he asked under which structural conditions the quotient procedure leads inevitably to the familiar bosonic or fermionic algebras. His starting point was to impose stability requirements on the ideal that generates the equivalence relation.
@@ III.1 Internal degeneracies and quantum grammar
We begin fixing a factorisation $H\cong\mathbb{C}^{d}\otimes K$ of the 1-particle space, with $E$ the external degrees of freedom and $K$ certain internal degeneracy space. Let $\{e_{i}\}_{i=1}^{d}$ be the standard basis of $\mathbb{C}^{d}$ and let $\{f_{\alpha}\}_{\alpha=1}^{m}$ be a basis of $K$. We define in this way generators
@@
The external index $i$ labels the external modes and transforms by a representation of $\mathrm{U}(d)$, while the internal index $\alpha$ labels hidden degrees of freedom and transforms in a way we do not control in principle. We will not fix the latter action hence; instead, it will be shaped by the quadratic relations.
@@ III.2 Order implies quadratic generation
The intuition behind the proof follows a physical line of reasoning. The transition from distinguishable to indistinguishable particles can be viewed as introducing relations among the various polynomials one can form from the chosen reference basis, which plays the role of an alphabet. Imposing that the many–particle space admits an ordered basis means precisely that, after applying the allowed relations, any monomial that does not respect the fixed order can be rewritten as a linear combination of ordered monomials.
@@ III.3 $\mathrm{U}(1)$ grading
The tensor algebra carries a natural $\mathrm{U}(1)$ action generated by the counting of particles: a phase $z\in\mathrm{U}(1)$ acts on $H$ by $X_{i\alpha}\mapsto zX_{i\alpha}$ and extends multiplicatively to $T(H)$. We will say that an homogeneous ideal $I$ is $\mathrm{U}(1)$–graded if it is invariant under this action. In that case the quotient $F$ is also $\mathrm{U}(1)$–graded, and the degree $n$ component transforms with weight $z^{n}$.
This grading expresses the compatibility of the statistics with local phase transformations that act diagonally in particle number. It is automatically satisfied for homogeneous quadratic ideals, so we take it as part of the basic setup rather than a separate constraint. The important thing is that assumption of an ordered basis implies that the quotient space carries a canonical decomposition.
@@ III.4 $\mathrm{U}(d)$ invariance
The first non–trivial constraint comes from covariance under external mode transformations. For this to be true for every state, the whole space of states should carry a representation of the unitary group on the external modes $U(E)$. For this to be true the ideal itself and especially the generating subspace should have a representation as well of such group. The external Hilbert space $E\cong\mathbb{C}^{d}$ carries the defining representation of $\mathrm{U}(d)$, and $H\cong E\otimes K$ inherits an action in which $\mathrm{U}(d)$ acts on $E$ and trivially on $K$. The tensor square decomposes as
@@ III.5 Quadratic relations and the generating ideal
where the tensors $K_{\mathrm{sym}}$ and $K_{\mathrm{ext}}$ encode the internal projectors. What we have now is that the generating space identify precisely those states that will vanish after the quotient. In that sense, these vectors we have just described are the ones that determine the generating relations of the algebra
@@
We can denote then corresponding quotient algebra by
@@ III.6 Yang–Baxter constraints and PBW property
All of this construction has been made over the assumption of the existence of an ordered basis after the quotient is performed. However, for the given projections and in consequence the generating relations just introduced to be compatible with such assumption we have to impose what we call the Poincaré–Birkhoff–Witt condition (PBW). In the language of quadratic algebras this amounts to requiring that the reduction of words using the quadratic relations is confluent [^1] A given word in a quadraticaly generated algebra can be reduced in different ways. The rewriting system given by the defining relations is confluent if no matter which sequence of allowed reductions you follow, you will end up with the same final form. A convenient formulation uses projectors rather than explicit relations.
@@ III.7 Single–mode partition function
An important feature of the construction that follows from our imposition of the ordered basis is the factorisation of the global Fock space into effective single–mode components. Given that we have that our resulting quotient space can be expressed as a tensor power of $U(1)$ representations as
@@
In the language of MedinaSanchezDakic2024 we know that the Hilbert-Poincare series of the algebra as defined here is in close relation with the partition function, once we exchange the variable $t$ with the Boltzmann factor $e^{\beta\,\epsilon}$. We can then finally state the main theorem:
@@ IV Creation–annihilation algebras and transfields
We now turn to the construction of the full creation–annihilation algebra corresponding to a given quadratic quotient of the tensor algebra consisten with transtatistics.
@@ IV.1 Annihilation algebra
This construction shows that the creation and annihilation algebras are related by orthogonality: the projector $P_{\mathrm{gen}}$ selects which quadratic combinations of creation operators vanish, while its orthogonal complement $P_{\mathrm{gen}}^{\perp}$ selects the ordered monomials that survive. Under the dual pairing these roles are reversed, so the annihilation relations mirror the creation relations with arrows reversed. In particular, the combinatorial structure governing annihilation is determined by the orthogonality between $R_{\mathrm{gen}}$ and $R_{\mathrm{gen}}^{\perp}$.
@@ IV.3 Vacuum two–point function and non–homogeneous term
Up to this point the construction determines only the homogeneous, quadratic part of the mixed relations. But this does not take in account so far the behaviour of vacuum.
@@ V An explicit finite–order example
To illustrate the abstract formalism we briefly describe a simple example with a finite maximal occupation number per mode.
@@ VI Conclusions and outlook
We have developed an operationally motivated generalisation of the transition from first to second quantization, without assuming symmetrisation principles as fundamental postulates. Building on the work presented in the previous chapter, we have provided an explicit algebraic formulation of families of transtatistics using quotients of tensor algebras over single-particle spaces. The assumption that an ordered basis of monomials survives the quotient then yields the operational locality condition, expressed as mode-by-mode independence of particle counting. Once this structural requirement is in place, we impose a U(d) symmetry on the accessible degrees of freedom, which fixes the external behaviour of the relations and supplies the representation-theoretic framework necessary for defining transtatistics.
From a physical perspective, this quotient-based construction offers a way of defining indistinguishability in operational terms, as the collapse of accessible information that could be used to differentiate between particles. Algebraically, the emergence of the Yang–Baxter equation as a necessary condition for statistical consistency links directly to the operational notion of locality: once states are defined solely through particle numbers, relative to additional information that is operationally inaccessible, the resulting space is quadratically generated, with the Yang–Baxter solutions ensuring invariance of the statistics under subsystem decompositions.
On the other hand, although we have defined the quadratic relations that determine the Fock space through the creation operators, the inner product on $\mathcal{H}$ allows us to introduce the annihilation operators in the standard way, namely as the adjoints of the creation operators. In the same way, we use this notion of orthogonality to formalize the creation and annihilation algebras. Moreover, this framework connects with existing work on parastatistics, which share the guiding idea of scrambling internal degrees of freedom that are operationally inaccessible, where we find the classical examples Green53 together with more recent developments WangHazzardProvided. In those cases the relations are introduced heuristically and by postulate, whereas here we obtain a complete characterization derived from basic operational assumptions about the quantum physics of many-particle systems.
@@ Appendix B Lemma 2
Finally, there is a natural $U(1)$-action on the quotient: an ordered monomial is multiplied by a phase whose argument depends on its degree. For $g\in U(1)$ we have $g\cdot w\coloneqq e^{in\theta}w$, for all $w$ of order $n$. In the same way, we have an action of $U(1)$ on $W^{\otimes d}$ similarly by adding up the degrees of the components on $w_{1}\otimes...\otimes w_{d}$ and then multiplying by a phase accordingly. Naturally this actions commute with respect to the isomorphism hence we have that
@@ 2. PBW Theorem
Direct application of the PBW theorem. ∎
@@ Appendix E Theorem 2
By construction: the ordered basis conditions implies the factorization property of the $U(1)$ character and additionally, the $U(d)$ invariance completes the conditions for the application of the partion theorem. The result follows. ∎

### v2

@@ Abstract
One of the traditional ways of introducing bosons and fermions is through creation–annihilation algebras. Historically, these have been associated with emission and absorption processes at the quantum level and are characteristic of the language of second quantization. In this work, we formulate the transition from first to second quantization by taking quotients of the state spaces of distinguishable particles, so that the resulting equivalence classes identify states that contain no information capable of distinguishing between particles, thereby generalising the usual symmetrisation procedure. Assuming that the resulting indistinguishable-particle space (i) admits an ordered basis compatible with how an observer may label the accessible modes, (ii) is invariant under unitary transformations of those modes, and (iii) supports particle counting as a mode-wise local operation, we derive a new class of creation–annihilation algebras. These algebras reproduce the partition functions of transtatistics, the maximal generalisations of bosons and fermions consistent with these operational principles.
@@ I Introduction
which describes distinguishable particles whose single–particle spaces are isomorphic. Indistinguishability is then imposed by projecting onto the symmetric or antisymmetric subspaces, yielding the usual occupation–number description in which only the number of particles per accessible mode is relevant. In this traditional formalism, the exclusivity of Bose and Fermi statistics is assumed rather than derived, and any alternative statistical behaviour is excluded a priori. Nonetheless, a variety of classical and recent works have exhibited consistent generalisations of statistics, including parastatistics Green53; Greenberg1964; HARTLE69; ryan1963; stolt1970; stoilova; WangHazzardProvided, quons Greenberg1991, Gentile statistics Gentile40, and several topological or combinatorial extensions dewitt; myrheim; Wilczek82; Read_2003; fredenhagen89; Surya_2004; Balachandran_2001; Baez06; vanenk2019; ZhangHanssonKivelson1989; Witten1989; Haldane08; Laughlin1983. These are typically introduced by ad hoc modifications of commutation relations, rather than through a structural analysis of indistinguishability.
In this work we develop a systematic treatment of the passage from distinguishable to indistinguishable particles based on operational constraints. We again take as starting point the algebra $T(H)$ whose elements represent labelled multi–particle states, following Segal’s tensor-algebra formulation of second quantization Segal1956a; Segal1956b; Segal1958. Indistinguishability is implemented by an equivalence relation that removes precisely the information operationally inaccessible for distinguishing one constituent from another. The quotient
@@
To connect this with experimental structure, we assume a factorization
@@ II.1 Operational indistinguishability and information loss
Consider a family of systems that, at the single–particle level, are described by the same Hilbert space $H$. An individual particle state is represented by a unit vector $\psi\in H$, and a collection of $n$ labelled particles is described by a vector in $H^{\otimes n}$. The labels encode individuality: the factor $H$ in each tensor position corresponds to a specific system, and permutations of the factors implement relabellings.
@@
In the standard approach, indistinguishability is represented by the action of the symmetric group $S_{n}$ on $H^{\otimes n}$, and the subspace of physical states is obtained by projecting onto the totally symmetric or totally antisymmetric subspace. This can be rephrased as choosing $I$ to be the ideal generated by the antisymmetric or symmetric components in degree two, respectively. The equivalence relation introduced in that case is the one that corresponds to the identification of states that only differ on a permutation of the labels. Depending on how the group $S_{n}$ is acting on the states, we will get either the symmetric or antisymmetric states.
@@ II.2 Fock space as a quotient algebra:
carries a natural $\mathbb{N}_{0}$–grading by tensor degree, which determines the number of particles. We assume that the ideal $I$ is homogeneous,
@@
In this way the transition from distinguishable to indistinguishable particles does not change the number of particles, as expected. Additionally we see that the degree–zero component $F_{0}$ is one–dimensional, spanned by the vacuum $|0\rangle$, and the degree–one component $F_{1}$ is naturally isomorphic to the one-particle space $H$.
@@ II.3 Segal’s theorem
The construction we employ was used by Segal, who asked under which structural conditions the quotient procedure leads inevitably to the familiar bosonic or fermionic algebras. His starting point was to impose stability requirements on the ideal that generates the equivalence relation.
@@ III.1 Internal degeneracies and quantum grammar
We begin by fixing a factorization $H\cong\mathbb{C}^{d}\otimes K$ of the 1-particle space, with $E$ the external degrees of freedom and $K$ certain internal degeneracy space. Let $\{e_{i}\}_{i=1}^{d}$ be the standard basis of $\mathbb{C}^{d}$ and let $\{f_{\alpha}\}_{\alpha=1}^{m}$ be a basis of $K$. We define in this way generators
@@
The external index $i$ labels the external modes and transforms by a representation of $\mathrm{U}(d)$, while the internal index $\alpha$ labels hidden degrees of freedom and transforms in a way we do not control in principle. We will not fix the latter action henceforth. Instead, it will be shaped by the quadratic relations.
@@ III.2 Order implies quadratic generation
The intuition behind the proof follows a physical line of reasoning. The transition from distinguishable to indistinguishable particles can be viewed as the introduction of relations among the various polynomials one can form from the chosen reference basis, which plays the role of an alphabet. Imposing that the many–particle space admits an ordered basis means precisely that, after applying the allowed relations, any monomial that does not respect the fixed order can be rewritten as a linear combination of ordered monomials.
@@ III.3 $\mathrm{U}(1)$ grading
The tensor algebra carries a natural $\mathrm{U}(1)$ action generated by the counting of particles: a phase $z\in\mathrm{U}(1)$ acts on $H$ by $X_{i\alpha}\mapsto zX_{i\alpha}$ and extends multiplicatively to $T(H)$. We will say that a homogeneous ideal $I$ is $\mathrm{U}(1)$–graded if it is invariant under this action. In that case the quotient $F$ is also $\mathrm{U}(1)$–graded, and the degree $n$ component transforms with weight $z^{n}$.
This grading expresses the compatibility of the statistics with local phase transformations that act diagonally in particle number. It is automatically satisfied for homogeneous quadratic ideals, so we take it as part of the basic setup rather than a separate constraint. The important thing is that the assumption of an ordered basis implies that the quotient space carries a canonical decomposition.
@@ III.4 $\mathrm{U}(d)$ invariance
The first non–trivial constraint comes from covariance under external mode transformations. For this to hold for every state, the whole space of states should carry a representation of the unitary group on the external modes $U(E)$. For this to be true the ideal itself and especially the generating subspace should have a a representation of this group as well. The external Hilbert space $E\cong\mathbb{C}^{d}$ carries the defining representation of $\mathrm{U}(d)$, and $H\cong E\otimes K$ inherits an action in which $\mathrm{U}(d)$ acts on $E$ and trivially on $K$. The tensor square decomposes as
@@ III.5 Quadratic relations and the generating ideal
where the tensors $K_{\mathrm{sym}}$ and $K_{\mathrm{ext}}$ encode the internal projectors. What we have now is that the generating space identifies precisely those states that will vanish after the quotient. In that sense, these vectors we have just described are the ones that determine the generating relations of the algebra
@@
We can then denote the corresponding quotient algebra by
@@ III.6 Yang–Baxter constraints and PBW property
All of this construction has been made over the assumption of the existence of an ordered basis after the quotient is performed. However, for the given projections and in consequence the generating relations just introduced to be compatible with this assumption we have to impose what we call the Poincaré–Birkhoff–Witt condition (PBW). In the language of quadratic algebras this amounts to requiring that the reduction of words using the quadratic relations is confluent [^1] A given word in a quadratically generated algebra can be reduced in different ways. The rewriting system given by the defining relations is confluent if no matter which sequence of allowed reductions you follow, you will end up with the same final form. A convenient formulation uses projectors rather than explicit relations.
@@ III.7 Single–mode partition function
An important feature of the construction that follows from our imposition of the ordered basis is the factorization of the global Fock space into effective single–mode components. Given that our resulting quotient space can be expressed as a tensor power of $U(1)$ representations as
@@
In the language of MedinaSanchezDakic2024 we know that the Hilbert-Poincaré series of the algebra as defined here is in closely related to the partition function, once we exchange the variable $t$ with the Boltzmann factor $e^{\beta\,\epsilon}$. We can then finally state the main theorem:
@@ IV Creation–annihilation algebras and transfields
We now turn to the construction of the full creation–annihilation algebra corresponding to a given quadratic quotient of the tensor algebra consistent with transtatistics.
@@ IV.1 Annihilation algebra
This construction shows that the creation–annihilation algebras are related by orthogonality: the projector $P_{\mathrm{gen}}$ selects which quadratic combinations of creation operators vanish, while its orthogonal complement $P_{\mathrm{gen}}^{\perp}$ selects the ordered monomials that survive. Under the dual pairing these roles are reversed, so the annihilation relations mirror the creation relations with arrows reversed. In particular, the combinatorial structure governing annihilation is determined by the orthogonality between $R_{\mathrm{gen}}$ and $R_{\mathrm{gen}}^{\perp}$.
@@ IV.3 Vacuum two–point function and non–homogeneous term
Up to this point the construction determines only the homogeneous, quadratic part of the mixed relations. But this does not yet take into account the behaviour of the vacuum.
@@ V An explicit finite–order example
To illustrate the abstract formalism we briefly describe a simple example with a finite maximal occupation number for each mode.
@@ VI Conclusions and outlook
We have developed an operationally motivated generalisation of the transition from first to second quantization, without assuming symmetrisation principles as fundamental postulates. Building on the work previously presented by the authors on the reconstruction of quantum statistics, we have provided an explicit algebraic formulation of families of transtatistics using quotients of tensor algebras over single-particle spaces. The assumption that an ordered basis of monomials survives the quotient then yields the operational locality condition, expressed as mode-by-mode independence of particle counting. Once this structural requirement is in place, we impose a U(d) symmetry on the accessible degrees of freedom, which fixes the external behaviour of the relations and supplies the representation-theoretic framework necessary for defining transtatistics.
From a physical perspective, this quotient-based construction offers a way of defining indistinguishability in operational terms, as the loss of accessible information that could be used to differentiate between particles. Algebraically, the emergence of the Yang–Baxter equation as a necessary condition for statistical consistency links directly to the operational notion of locality: once states are defined solely through particle numbers, relative to additional information that is operationally inaccessible, the resulting space is quadratically generated, with the Yang–Baxter solutions ensuring invariance of the statistics under subsystem decompositions.
On the other hand, although we have defined the quadratic relations that determine the Fock space through the creation operators, the inner product on $\mathcal{H}$ allows us to introduce the annihilation operators in the standard way, namely as the adjoints of the creation operators. In the same way, we use this notion of orthogonality to formalize the creation–annihilation algebras. Moreover, this framework connects with existing work on parastatistics, which share the guiding idea of scrambling internal degrees of freedom that are operationally inaccessible, where we find the classical examples Green53 together with more recent developments WangHazzardProvided. In those cases the relations are introduced heuristically and by postulate, whereas here we obtain a complete characterization derived from basic operational assumptions about the quantum physics of many-particle systems.
@@ Appendix B Lemma 2
Finally, there is a natural $U(1)$-action on the quotient: an ordered monomial is multiplied by a phase whose argument depends on its degree. For $g\in U(1)$ we have $g\cdot w\coloneqq e^{in\theta}w$, for all $w$ of order $n$. In the same way, we have an action of $U(1)$ on $W^{\otimes d}$ similarly by adding up the degrees of the components on $w_{1}\otimes...\otimes w_{d}$ and then multiplying by a phase accordingly. Naturally, these actions commute with respect to the isomorphism hence we have that
@@ 2. PBW Theorem
The result follows by direct application of the PBW theorem. ∎
@@ Appendix E Theorem 2
By construction: the ordered basis conditions implies the factorization property of the $U(1)$ character and additionally, the $U(d)$ invariance completes the conditions for the application of the partition theorem. The result follows. ∎